Seecompany is evaluating the replacement of an old machine with a new one. Last year, th cd a consultant to conduct a feasibilit above details.

1. What is the initial investment CF0 ?

2. What is the CF1 ( The Cash Flow to be used in NPV calculations)?

3. What is the CF4 ( The Cash Flow to be used in NPV calculations)?

4. What is the NPV of the project?
Answer #1

Answer 1:

Old machine:

Old Machine: Depreciation Schedule Year Depreciation rate Depreciation Accumulated depreciation $600,000 $1,560,000 $2,136,00

Book Value now = $1,440,000

Sale value = $800,000

Tax on Loss = (1440000 - 800000) * 40% = $256,000

Sale value net of tax =800000 + 256000 = $1,056,000

Workings capital:

Increase in working capital = Increase in inventory - Increase in accounts payable = 800000 - 200000 = $600,000

Initial investment CF0:

Initial investment CF0 = New Machine cost - sale value of old machine net of tax + Increase in working capital

= 2500000 - 1056000 + 600000

= $2,044,000

Initial investment CF0 = $2,044,000

Answer 2 and 3:

CF1 = $589,600

CF4 = $1,230,880

Workings:

Year New Machine Sale value of old machine net of tax Increase in working capital 0 $2,500,000 $1,056,000 ($600,000 2 4 $600,

Answer 4:

NPV of the project = $293,433.26

Workings:

Year New Machine Sale value of old machine net of tax Increase in working capital 0 $2,500,000 $1,056,000 ($600,000 2 4 $600,

---------------------------------------------------------------------------------------------------------------------------------------------------------------

Table added (please refer comment):

Row # | Year New Machine Sale value of old machine net of tax Increase in working capital 0 $2,500,000 $1,056,000 ($600,000 4